# Transformer-based RGB-T Tracking with Channel and Spatial Feature Fusion

### v1

@@ IV-E1 Heatmap
We explore how CSTN utilizes RGB and TIR modalities to improve model performance. As shown in Figure 7, in the bikeboyright sequence which is a typical daytime traffic scenario, our proposed CSTNet effectively uses RGB and TIR modalities to better predict the target state. In the biketurnright sequence which is a typical nighttime scenario, although CSTNet is interfered with in RGB feature, the model still makes a correct prediction in fused features under the reference of TIR feature. In the whitebikebelow and yellowgirl118 sequences, CSTNet improves target state prediction in high illumination and partial occlusion challenges by complementing the features of RGB and TIR modalities. CSTNet effectively achieves cross-modal feature interaction.
In addition, we compare the cross-modal feature interaction performance of CSTNet and other Transformer-based methods (TBSI [2] and BAT [15]) in four typical sequences of the LasHeR [28] dataset, as shown in Figure 8. Our proposed CSTNet achieves direct interaction between RGB and TIR original semantics through channel and spatial fusion of multimodal features. In the basketboy, bikefromlight, and motocomeonlight sequences, CSTNet achieves richer utilization of cross-modal appearance features. In the blkstandboy sequence, CSTNet has better discrimination for similar appearance. Overall, CSTNet demonstrates the benefits of direct cross-modal feature interaction on model performance.
IV-E2 Tracking Results
@@
We select four representative sequences from the Lasher dataset to show some of the results of CSTNet and other previous state-of-the-art RGB-T trackers. As shown in Figure 9, in the rightdarksingleman sequence, when other trackers briefly lose the target due to low light and camera shake, our proposed CSTNet still tracks the target stably, demonstrating the better robustness of CSTNet. In the ab_blkskirtgirl sequence, when the target is frequently occluded, other trackers usually lose the target briefly, while our CSTNet does not lose the target. In the bikefromlight sequence, when the scene is highly illuminated, our CSTNet accurately distinguishes the target from surrounding clutter, resulting in better performance. In the boytakingplate2left sequence, our CSTNet better integrates the RGB and TIR features of the target, resulting in better performance in partial occlusion and similar appearance challenges. Overall, our proposed CSTNet has stronger discriminative ability and better tracking performance.
@@ V Conclusion
In this paper, we present a new RGB-T tracker CSTNet. CSTNet explores the advantages of using channel and spatial feature fusion to achieve direct interaction between RGB and TIR features. CSTNet contains two novel cross-modal feature interaction modules, CFM and SFM. The CFM achieves joint channel enhancement of RGB and TIR features, joint multi-level modeling, and global integration of fused features and original features. The SFM achieves spatial mutual modeling of RGB and TIR features and joint spatial and channel integration of multimodal features. Comprehensive experiments show that the direct interaction between RGB and TIR features improves the performance of the model, and our proposed CSTNet achieves state-of-the-art performance on three challenging RGB-T tracking benchmarks.

### v2

@@ IV-E3 Ablation of SFM
IV-E4 Inserting Layers of CFM and SFM
In addition, we conduct ablation experiments on different insert layers of our proposed CFM and SFM modules. The results are reported in Table VIII. Inserting the CFM and SFM modules in the 4th layer of the ViT improves the PR by 0.4% and the SR by 0.3%. Inserting the modules in the 7th and 10th layers of ViT further improves the performance of the tracker. Overall, inserting our CFM and SFM modules into these three layers effectively achieves feature fusion at different stages and improves model performance.
IV-E5 Ablation of pre-trained models
We explore the impact of different pre-trained models on our proposed CSTNet in terms of performance. CSTNet without pretraining and CSTNet with ImageNet pretraining are unable to converge in 20 epochs, therefore we increase their learning rate by 10 times. The results are shown in Table IX, CSTNet with RGB-T pretraining achieves the best performance. It outperforms CSTNet with SOT pretraining by 1.8% for PR and 1.4% for SR. The RGB-T pre-training model achieves better parameter initialization of our model, reducing our training epochs to one-third compared to ViPT [1], TATrack [17], and STMT [41] and BAT [16]. This shows that using the RGB-T pre-trained model can further improve the performance of the tracker.
The use of pre-trained RGB-T models is not intended to gain an inappropriate performance advantage; rather, we hope to promote the development of RGB-T trackers through a better pre-trained model, as in the case of [23].
IV-F Visualization
IV-F1 Heatmap
We explore how CSTN utilizes RGB and TIR modalities to improve model performance. As shown in Figure 8, in the bikeboyright sequence, which is a typical daytime traffic scenario, our proposed CSTNet effectively uses RGB and TIR modalities to better predict the target state. In the biketurnright sequence which is a typical nighttime scenario, although CSTNet is disturbed in RGB feature, the model still makes a correct prediction in fused features under the reference of TIR feature. In the whitebikebelow and yellowgirl118 sequences, CSTNet improves target state prediction in high illumination and partial occlusion challenges by complementing the features of RGB and TIR modalities. CSTNet effectively achieves cross-modal feature interaction.
In addition, we compare the cross-modal feature interaction performance of CSTNet and other Transformer-based methods (TBSI [2] and BAT [16]) on four typical sequences of the LasHeR [31] dataset, as shown in Figure 9. Our proposed CSTNet achieves direct interaction between RGB and TIR original semantics through channel and spatial fusion of multimodal features. In the basketboy, bikefromlight, and motocomeonlight sequences, CSTNet achieves richer utilization of cross-modal appearance features. In the blkstandboy sequence, CSTNet has better discrimination for similar appearance. Overall, CSTNet demonstrates the benefits of direct cross-modal feature interaction on model performance.
IV-F2 Tracking Results
@@
We select four representative sequences from the Lasher dataset to show some of the results of CSTNet and other previous state-of-the-art RGB-T trackers. As shown in Figure 10, in the rightdarksingleman sequence, when other trackers briefly lose the target due to low light and camera shake, our proposed CSTNet still tracks the target stably, demonstrating the better robustness of CSTNet. In the ab_blkskirtgirl sequence, when the target is frequently occluded, other trackers usually lose the target briefly, while our CSTNet does not lose the target. In the bikefromlight sequence, when the scene is highly illuminated, our CSTNet accurately distinguishes the target from the surrounding clutter, resulting in better performance. In the boytakingplate2left sequence, our CSTNet better integrates the RGB and TIR features of the target, resulting in better performance in partial occlusion and similar appearance challenges. Overall, our proposed CSTNet has stronger discriminative ability and better tracking performance.
@@ V Conclusion
In this paper, we present a new RGB-T tracker CSTNet. CSTNet explores the advantages of using channel and spatial feature fusion to achieve direct interaction between RGB and TIR features. CSTNet contains two novel cross-modal feature interaction modules, CFM and SFM. The CFM achieves joint channel enhancement of RGB and TIR features, joint multi-level modeling, and global integration of fused and original features. The SFM achieves spatial mutual modeling of RGB and TIR features and joint spatial and channel integration of multimodal features. Comprehensive experiments show that the direct interaction between RGB and TIR features improves the performance of the model, and our proposed CSTNet achieves state-of-the-art performance on three challenging RGB-T tracking benchmarks. We also provide a small version, CSTNet-small. It achieves high performance RGB-T tracking using the model without the CFM and SFM modules and CSTNet weights as the pre-training weights. We evaluate the parameters, flops, and speed of CSTNet and CSTNet-small, and deploy them to a embedded computing device. Comprehensive experiments demonstrate that our models meet the real-time requirement of practical applications.